HOW TOs

From SBS wiki
Jump to: navigation, search
RC phone: 757-876-1787
Crew chief: x7045
Operator 1: x7046
Operator 2: x7047
Subsystem Expert Contact Info
GMn Contact Information
Shift Schedule
Shift checklist
Shift Checklist How Tos
Hall A Phone Book
SBS Shifters Wiki Info
Documentation
HOW TOs for shift crew
Expert Tools
Daily RC Meetings
<< SBS Main

This is the place for task-specific instructions like rebooting ROC XX, starting the High Voltage GUI or bringing up the Alarm Handler.

Contents

Beam Recovery Procedure for GEN-RP/K_LL

  • Initially MCC will want to send tune beam into the hall to establish a good beam orbit by following this accelerator procedure: [1]. Before any beam is sent into the hall make sure the target ladder is in the "NO TARGET" position meaning all targets are out of beam. Also make sure the Moller target ladder is out of beam. All SBS and BB detectors should be off. SBS and BB magnets should be off as well. Also check, if the Moller Quads are OFF or not (field integral should be zero during GEn-RP/K_LL as in [2]). If ON, ask MCC to degauss the Moller Quads and turn off power supplies.


  • As part of the accelerator procedure mentioned above the intrinsic beam spot size (unrastered beam spot size) will be checked by performing harp scans with IHA1H04A and IHA1H04B. MCC will do harp scans with IHA1H04A and IHA1H04B with 5 muA TUNE beam with the target ladder in the "NO TARGET" position and raster OFF.


  • Once MCC posts the harp scans to the elog check the beam width at the target in X and Y using instruction below To get the beam width from harp scans. If the beam width at the target is between 200 and 300 microns move on to the next step. If not, ask MCC to work on the beam profile. Make a HALOG with the final results (example here [3]).


  • MCC/Hall A will then want to assess the level of beam scraping, if any, on the differential pump region small aperture in the downstream beam line right after the target [4], [5].


  • At this stage MCC will need to do a functional test of the ion chambers following this procedure:[6]. I hope the downstream IC that's under the platform is easy enough to trip during this procedure so we don't waste time.


  • Once the IC functional test is done we can start working on centering the beam on the Carbon hole. There are 2 Carbon foils with holes at the center placed on different holders: one has a 2 mm hole while the other a 5 mm hole [7].
    • Turn ON the BigBite magnet and the BigBite detectors. If BB GEMs are needed, consult with the RC and GEM expert.
    • Select the GEnII_NoSBS coda config and set PS1 to 0 while all other triggers are off in the PS gui.
    • Ask MCC to set the x,y beam positions at IPM1H04A and E to zero in the orbit locks gui. It is a good idea to check that the XSOF and YSOF have been returned to their original values [8].
    • Request 4x4 mm rastered CW beam at 5 muA and take a run on the 2 mm C-hole target. Replay the run and see if you can identify the 2 mm hole by checking the raster plots. Keep in mind that we only use the first pair of raster coils to raster the beam for GEN_RP/K_LL.
    • Adjust the beam positions until the 2 mm C-hole is centered in the 4 mm raster pattern.
    • Once the Carbon hole is centered into the raster pattern ask MCC to set the beam position offsets at IPM1H04A and E in the orbit lock gui to the just found alignment numbers.


  • Reduce the raster size until the 2 mm C-hole is concentric within the raster pattern. This will give you the 2x2 mm raster needed for production.


  • This paragraph is written by Ciprian, feel free to contact him for questions: "The raster size can be x-checked with the harps: see https://logbooks.jlab.org/entry/3725647 for example. Have MCC turn on the raster, send up to 5uA tune beam and do a swipe of the 1H04A and 1H04B harps. The fits will look bad (see the example) but that is ok - make sure the MCC operator doesn't run multiple harp swipes because their fit is not converging. Take the width and height of the raster from the harp swipe (by eye) and put them in the same code that you used for the beam width to get the value at the target."

(!! Old !!) Beam Recovery Procedure (Followed during GEn for Glass target)

If beam has been down for over 30 minutes move target to "No Target"

This quick procedure applies to coming back from a shorter period of no beam. If coming back from machine tuning or any work that might affect the beam profile, the full procedure below should be executed.

  1. Move to No Target position
  2. Verify the beam setup
    Target BPM locks at (0,0)
    For He3 production target (and no target position) BPM 4A XSOF, YSOF = (-0.10, 1.20) 4E XSOF, YSOF = (-0.32, 1.07)
    Raster size: Vertical: 1.75V Horizontal: 1.50V.
    Check the beam position orbit locks are engaged with correct values
    Check Hall A BPM Target Protection is Enabled
    Check Hall A Current Ramp Protection is Enabled
  3. Make a log entry like this: https://logbooks.jlab.org/entry/4110171
  4. Ask MCC to send 5uA beam with the setup above
  5. Monitor the target BPM positions. Once MCC can deliver stable beam with No Target, return to PolHe3 cell
  6. Do not change the setup other than moving the target to He3. NEW Check the voltage encoder readback is approximately 7.08 +/- 0.01 in target motion GUI</s>
  7. Ask MCC to send 5uA CW beam and monitor all above, and additionally IC + trigger rate

If beam has been down due to machine tuning

  1. Move to the "no target" position (detectors HVs OFF: especially GEMs should be off) before MCC starts to tune beam to the hall [consult with RC or sbs beamline experts to verify the procedure and numbers below]
  2. If the raster is ON, call MCC and ask for the raster to be turned OFF
  3. Request harp scans at IHA1H04A and IHA1H04B with 5 muA of TUNE beam with the target ladder in the "NO TARGET" position and raster OFF. Check if the Moller Quads are OFF or not (Field integral should be zero during GEnII as in [9]). If ON, ask MCC to degauss the Moller Quads and turn off Power supply
  4. Check the beam width at the target in X and Y using the instruction below To get the beam width from harp scans. If the beam width at the target is between 200 and 500 microns move on to the next step. If not, ask MCC to work on the beam profile. Make a HALOG with final result (example here [10])
  5. Once the beam width at the target in X and Y is between 200 and 500 microns, ask MCC to turn the raster ON and to set the Vertical amplitude 1.75 V and horizontal amplitude 1.50 V. We are using one pair of raster controlled by single generator 1, so Raster A should be unmasked and Raster B should be masked as in the example [11] . Please check that the amplitudes are as stated above in both the sending and readback fields for Generator 1. Also check that the Ix and Iy currents are close to what's shown in the picture posted at the link above. Also look towards the scope in the electronics room, you should see a nice circular pattern of the raster on the scope.
  6. Now that you are sure that the raster is set up correctly you can call MCC and tell them that you want to move the target ladder to the Carbon hole target to check both that the raster size looks like what we had before and also the beam centering.
  7. You are now on the Carbon hole target. Make sure MCC sets the Target BPM XSOF and YSOF
    a) The IPM1H0A,E are set to (0,0), and their XSOFs, YSOFs should be
    b) IPM1H04A (XPOS, YPOS) = (-0.10, 1.20)
    c) IPM1H04E (XSOF, YSOF) = (-0.32, 1.07)
    d) Make sure the target orbit lock is ON.
    e) The raster is already ON from the previous step but check anyway.
  8. Take a CODA run Turn ON all detector HV (don't forget to turn ON specially B0igbite shower, preshower and BB GEMs). Use prescale ps1=0 for this run to accumulate statistics fast, and take a run with ~10 uA CW and the production BBCal thresholds. [note: start a run only after BPMs come closer to nominal values, as beam positions might be off at the beginning]. Please keep eyes on IC rates during step 8 above while ramping current, comparing with [ Please keep eyes on IC rates during step 8 above while ramping current, comparing with [12] [13] ][14] [15] Once you have 200k events, go to the next step while continuing to take data.
  9. Analyze 50k and 200k events
  10. Call RC or Hall A beamline expert to get approval. If the hole is no clearly seen, you can run this script[16]
  11. Now you determined that: the intrinsic beam width is acceptable, the raster is rastering the beam as expected for data taking on glass cells and that the rastered beam is centered on the Carbon hole targets. You can move on to whatever the run plan requires. If moving to a glass cell first establish the beam position using the appropriate x/ysofs. Also check that the orbit locks, ramp protection, and target bpm protection are all on. Make a HALOG like this [17]
  12. Before requesting beam, be sure to read below: What to monitor while running on glass cells [18]
    Please follow the beam positions carefully as well as the Ion Chamber readings. You can load the typical stripchart I am using by right-clicking on any stripchart and loading the config from SBS/simona/beam/instructions.
    Also make sure the ramp protection is on, the target orbit lock is on and that the raster is on. To verify that the ramp protection is on type "jmenu" in any aonl or adaq terminal. Once the jmenu gui pops up type "ramp" in the search field and click enter. Select from the new menu "Hall A ramp protection" and make sure everything is green like in this example [19]

To get the beam width projected at the target from harp scans

  1. ssh a-onl@aonl1
  2. gosbs
  3. analyzer beam_width.C


What to monitor during production running on glass cells

1. Raster always ON when running on glass cell targets. Raster size is 5.2mm, V: 1.75 V H: 1.50 V. Cross check with the values on the white board.

2. Make sure the correctors that maintain the beam position close to the the ideals have changing Bdls meaning the target orbit lock is running.

3. There might be instances when although the target orbit lock is running the beam positions slowly drift from the ideal due to orbit instability in the machine: [20], [21]. In this case call MCC and ask them to investigate. Also call the Hall A beamline expert.

4. Ion chamber readings. Elevated IC readings compared to a baseline established by the beamline expert are indicative of scraping. The baseline for the target ion chambers for running on He3 at 3 pass is here: [22][23][24]

5. BBcal trigger rates. The BBcal trigger rate for a threshold of 57 mV as a function of current when running on He3 is posted here: [25][26] [27] [28] [29] and here after target parameter changes [30]. Rates from H2 running: [31]

6. Raster plots to ensure there's no scraping. A case of mild brushing of C foils holder by the beam envelop is shown here [32]. The way you want this to look it's like here [33]

7. "Hall A ramp protection" always ON, make sure everything is green like in this example [34]

8. NEW Check the voltage encoder readback is approximately 7.08 +/- 0.01 in target motion GUI

9. If beam has been gone for more than 30 minutes, always go to no target

CODA and DAQ

[Stopping and starting a run]

To monitor the disk usage, use the script "diskmon" on adaq@adaq1. This will show the fractional disk usage on each of the three disks, updating every two minutes. If the disk fraction is over 95% for the present output disk, contact Alexandre to have him change output disk.

Checking total event number for past run: You can check the number of events from a past run in a terminal:

ssh -Y a-onl@aonl1
cd gmn
./get_final_events_from_dalma.sh *run number*

The DAQ will not start or is crashing? LOOK HERE

Fixing the SBS DAQ

Running the parity DAQ

The parity DAQ will be running in a VNC session on adaq@adaq3.

  1. Log in to adaq@adaq3
  2. Execute "vncviewer -Shared :40"
  3. The parity DAQ should already be open.
    • If you need to start the parity DAQ
      1. Open a terminal window in the VNC and log in to apar@adaq3
      2. Execute "kcoda". Maybe do it again just to be sure.
      3. Execute "startcoda". This will launch a bunch of xterms and the RunControl window
      4. Within the RunControl window, select "Configurations" and then "Cool" on the top menu bar to bring up the run type list. Choose the "CntHouse" run type.
      5. Then do the "configure" and "download" steps like on the SBS DAQ
  4. We are running the parity DAQ in auto-mode with up to 50 runs of one hour each. Use the "start" button on RunControl instead of the "Prestart"/"Go" sequence.
    This is indicated here: ParityDAQ runControl start.png
    • If you need to change the number of runs remaining in the auto-mode sequence:
      ParityDAQ runControl scheduler.png
      1. Select "Options" on the top menu bar, then the sub-menu items "Scheduler"/"Program...". A window will pop-up, with sliders and text boxes for the number of runs in the sequence and the event and time limits per run. Type "50" in the number of runs and hit enter, and similarly type 60 as the time limit in minutes and hit enter. The press the "Set" button on the menu.
      2. If you had already started a run with "Start" the RunControl title bar should show that we're in auto-mode with the specified time limit and number of scheduled runs. This will update when you set new values as described above.
      3. If the RunControl title bar does not show that we're in auto-mode with the time limit and number of runs you've set, then stop the run and restart a new run with the "Start" button. The title bar should now show that we're in auto-mode.

What to do when the IHWP state is changed

When the IHWP is taken out or inserted, we need to run a script to change the injector control voltages.

  1. Be sure that both the SBS DAQ and parity DAQ runs have been stopped before changing the IHWP
  2. Log in to apar@adaq3. If you have the parity DAQ VNC open, one of the terminals will already be there.
    1. "cd ~/feedback2022"
    2. "./FlipPCValuesSBS".
      • This will ask you to verify the DAQ runs have been stopped, and it will then tell you the current IHWP state to verify that the IHWP state was changed. If you answer yes to both questions, it will then update the injector control voltages to be appropriate to whichever state the IHWP is in at that moment.
      • The flipper script will ask if you want it to post a log entry. Please say yes.
    3. After you run the flipper script, restart the parity DAQ using the "START RUN" button (the two arrows pointing right), not by pressing "Prestart/Go". If you use "Prestart/Go", you will not start the automatic mode and we will miss data!
    4. If you want to check the settings at any time, you can run "./checkPCvalues", also from the feedback2022 directory on apar@adaq3.

Expert instructions for parity DAQ analysis and working with feedback

To analyze a parity run

  1. ssh apar@adaq3
  2. source ~/PREX/setup_japan_SBS.tcsh
  3. go_feedback: this puts you into ~/PREX/japan_feedback-SBS/
  4. ./build/qwparity --config sbs_CntHouse.conf -r <parity runnumber>: this will produce a rootfile at ./japanOutput/sbs_CntHouse_<runnum>.root, also avalable as $QW_ROOTFILES/sbs_CntHouse_<runnum>.root
  5. root $QW_ROOTFILES/sbs_CntHouse_<runnum>.root
    1. mul->Draw("yield_bcm_dg_ds:Entry$","ErrorFlag==0"): this shows the beam current as a function of time
    2. mul->Draw("asym_bcm_dg_ds*1e6","ErrorFlag==0"): this is the BCM asymmetry in ppm for that parity run

To run the charge asymmetry feedback

  1. ssh apar@adaq3
  2. source ~/PREX/setup_japan_SBS.tcsh
  3. go_feedback: this puts you into ~/PREX/japan_feedback-SBS/
  4. feedback_ana: this starts the feedback process and stores a copy of the log outputs in "LogFiles"

To update the PITA1/PITA2 settings we use in the flipper script

  1. ssh apar@adaq3
  2. cd ~/feedback2022
  • Now you will need to update the file FlipPCValuesSBS to change one or both of these values (the -2700 for OUT and -2240 for in were determined in early October 2022):
Ref_PITA_OUT=-2700
Ref_PITA_IN=-2240
  • You will also need to update the file checkPCvalues to change the ideal PITA values and the IN/OUT ranges:
Ref_PITA_OUT=-2700
Ref_PITA_IN=-2240
Ref_PITA_OUT_HI=-2450
Ref_PITA_OUT_LO=-2950
Ref_PITA_IN_HI=-2000
Ref_PITA_IN_LO=-2500

To update the CODA scheduler parameters

  1. In CODA RunControl, select the "Options" menu on the upper menu bar, then the "Scheduler/Program" submenu
  2. Set the number of runs and time limit as desired, then push the "Set" button

Some very special commands that we might later use to make the parityDAQ/feedback more automatic

  • plask -rt CntHouse -schedulerStatus: Queries the state of the CODA scheduler
  • plcmd -rt CntHouse -startrun: Starts a coda run
  • plcmd -rt CntHouse -end: Ends a coda run; however, if the scheduler is active it will start the next run in the schedule

HV Controls

Most detector HV is controlled through the main gui. To access this:

  • Log into aslow@adaqsc
  • execute go_hv

To open the GEn-RP HV GUI for the recoil polarimeter hodoscope and active analyzer

  • Log into aslow@adaqsc
  • execute cd slowc/
  • execute ./hvs LEFT
  • Currently slots 9 - 12 are for the 48 hodoscope PMTs

This will open the HV GUI which will have buttons for each of the detectors. Shift workers can only turn HV on or off. Never change the HV numbers.

Shift Checklist Guide

Shift Checklist HowTos

Setting prescales and trigger

The Trigger Supervisor (TS) makes all triggers available at any time. The triggers that actually are included in the datastream are controlled by the prescale factors set at run time. To set the prescales, from any terminal on an hadesk computer log in to either adaq1 or adaq2 as follows:

ssh -Y sbs-onl@adaq1
prescale

Trigger supervisor (TS) inputs and corresponding prescales are shown in the table below. Note that PSX=-1 turns off trigger X. Otherwise, the prescale factor is approximately 2^PS (or really 1+2^(PS-1) )

Trigger # Prescale # Signal
T1 PS1 BBCAL
T2 PS2 HCAL
T3 PS3 BB && HCAL Coinc
T4 PS4 LHRS
T5 PS5 GRINCH Pulser
T6 PS6 HCAL LED
T7 PS7 BBCAL Lo (cosmics)
T8 PS8

For data-taking, the normal prescale settings are: PS1=0 PS6=0 all other PSn= -1, if GRINCH data is not needed, or PS1=0 PS6=0 and PS6=4

PS1=0 allows the BB-shower to trigger the DAQ.

PS6=0 allows a DVCS pulser (normally set to 10 Hz), which triggers the HCal LED-calibration system, to trigger the DAQ.

PS5=4 reduces the GRINCH LED pulser from 100 Hz to roughly 10 Hz.

If you suspect the DVCS pulser for the HCal LEDs is not set to 10 Hz, (eg. beam-off rate is not 30 Hz or 40 Hz, as expected, from the 20 Hz dead-time monitor(EDTM) and the 10 Hz of HCal LED events and 10 Hz of GRINCH LED, if enabled.) then, from any window logged on as adaq or sbs-onl:

ssh -Y daq@enpcamsonne
cd test_fadc
python3 DVCS_Pulser_Control_GUI.py

This will launch a GUI. Your terminal screen should indicate the last setting of the rate. If it's not 10 Hz, or if you don't trust it. Click "Load DVCS Pulser Library" then click "Initalize DVCS Pulser" then click "Enable Output Channels". This will stop the pulser. Now click on the rate you want, usually 10 Hz.

EDTM instructions at https://hallaweb.jlab.org/wiki/index.php/How_to_Set_EDTM_Frequency_with_the_DVCS_Pulser

Analyzing the data

Instructions for shift crew

  • 1) Login to a-onl@aonl1 and execute: (use this in aonl1 machine)
gosbs
./run_genrp_large_aonl1.sh runNum nEvents(in thousand) nSegments
  • 2) runNum is the run number you want to analyze and nEvents should be in terms of thousands (eg for 100k events to replay, use 100). If number of events in not provided, it will analyze 50k events from this run by default. nSegments should normally be 5 and divides the job into segments for faster processing.
  • 3) Please review all produced plots of each GUI, and click on "Exit GUI", this will pop the next GUI.
    • Example plots and instructions for understanding what to look for in HCal summary plots are here. Note: Compare the example plots with your 50k replays, and use the written explanations if there are discrepancies.
  • 4) After you're reviewed all plots, they will be saved as PDFs. YOU will be prompted in the terminal to answer (y/n) to post the plots in the HALOG. Ordinarily you will always answer yes ("y"), unless something went wrong with the replay.
  • 5) Each shift should try to have 1 (and only 1) full replay running at all times. It takes 4-5 hours so you will only be able to complete one or two full replays per shift. To run,
   gosbs
  ./full_replay_genrp.sh runNum 

How to get the run charge

  • Login to a-onl@aonlX (X=1, 2, 4) and execute:
gosbs
getCharge.sh runNum
  • This will print out the charge in Coulombs for that run. Please write it down on the run list spreadsheet.


Monitor and Control GUIs

Most beam/accelerator related GUIs are accessible via jmenu. To bring up jmenu log in from a Counting House computer as adaq@adaq1 then type "jmenu" at the command prompt. Alternately you can access jmenu via the OPS account "hacuser", for example by running ssh -X hacuser@hlal00.acc.jlab.org 'jmenu' which will require a password unless you are logged in as adaq on one of the adaq machines. The jmenu search feature is pretty good at finding relevant GUIs. Use "halla" in your search to include only variables relevant to Hall A.

Opening Specific jmenu GUIs

Bringing Up Main GUIs From Scratch

To bring the main GUIs we want to monitor on the TV in the counting house (hatv1), in a terminal (on any computer) type:

ssh aslow@adaqsc
start_all_vnc

This will start 6 VNC servers on adaqsc with addresses :1 thru :6. Each server is dedicated to a particular slow controls GUI and its geometry is optimized for that GUI. If any servers are already running, they will be terminated and restarted. Once the servers have started, you can log out from adaqsc.

Next, we need to connect to the VNC servers from hatv1. In a terminal on hatv1, make sure that you are logged in as aslow. (If you are not, sign in by typing ssh -Y aslow@hatv1. Be sure to do this on hatv1, so that your display is the TV screens.) In this terminal, type vncviewer &. This will bring up a window with icons for all VNC servers on adaqsc (and maybe a few others), labeled "# - GUI Name", where # = 1...6. Click each icon to open a vncviewer window for that GUI. Since we are starting from scratch, you should see only a blank desktop in that VNC session. Bring up a terminal on that desktop and type start_tool_screens [option], where [option] is one of

  • GeneralTools = (accelerator overview)
  • HallATools = (LHRS, beam line, gas systems, etc)
  • Raster
  • BBctrl = (BigBite magnet current, voltage, etc)
  • SBSctrl = (Super BigBite magnet current, magnetic field, and corrector magnets)
  • Beam = (beam energy, trajectories, positions)

and should match the name of the VNC window where you start it. Note: the options are: [GeneralTools, HallATools, Raster, BBctrl, SBSctrl, Beam]. You'll have to run this command 6 times, once in each VNC session, and each time with a different option, the one matching the server number you connected to.

Once all the GUIs are up in the VNC sessions, resize the VNC screens to your liking on hatv1.

An astute observer might notice that this elaborate procedure is essentially a workaround for the lack of scaling support in EPICS GUIs. When resizing an EPICS GUI window, its widgets remain fixed, and only the window canvas will change. The TVs connected to hatv1 are 4K devices on which the Accelerator GUIs, written for 1990s-era low resolution screens, look veritably microscopic. To enlarge the GUIs to a readable size, we take advantage of the fact that the VNC viewer application we are using supports continuous, arbitrary scaling of the entire desktop and thus of all windows inside it. Only the RealVNC viewer has this support, not the TigerVNC viewer.

Troubleshooting

If the GUIs disappear and you see completely black screens within the VNC sessions, even after restarting, the following procedure is usually successful: As aslow@adaqsc, terminate the defunct VNC sessions: vncserver -kill :#, where #=1...6. Then kill all runaway gnome-keyring-daemon processes, if necessary with kill -9. Next, still on adaqsc, do sudo systemctl daemon-reload, which will unregister the stale desktop sessions from systemd and allow you to restart the GUIs from scratch, as described above. aslow is allowed to run this command as sudo.

Bringing Up Scaler GUI

To bring up the Hall A Scaler GUI, which displays the scalers from the BigBite SHower Trigger sums, beam line, and trigger, ssh to a-onl@adaqX (X = 1, 2, 3, etc) type:

goxscaler

Then type ./xscaler SBS or ./xscaler Left.

Plotting EPICS variables with MyaPlot or LivePlot

For more information see Strip Chart

  1. From the jmenu GUI select Plots>>MyaPlot or Plots>>LivePlot
  2. Type into the plot the name of the EPICS variable(s) you wish to view. If you can't recall the name, you can always search jmenu for a GUI where such a variable might be displayed. For example, the Hall A Moller target position can be obtained from jmenu by searching for "halla moller target". Clicking the middle mouse button on the control or display variable located on the GUI copies the variable name into the buffer. Middle clicking again on the MyaViewer where the EPICS variable is entered will paste it into the plot.

Procedure for Ramping the BigBite Magnet

The steps to ramp the BigBite magnet are as follows. The nominal current is 750 A (note that this has changed for kinematic 3 and 4 from 700A). The nominal polarity is NEGATIVE.

  1. Obtain permission from the Run Coordinator to ramp the BB magnetic field.
  2. Make a post in the HALOG indicating that the BigBite magnetic field will be ramped to the desired current setting.
  3. Bring up the BigBite operational GUI. On any adaq machine, when logged in as aslow or adaq, type start_tool_screens BBctrl
  4. If this command does't bring up the BB magnet GUI, you can access it via the jmenu as well. Open a jmenu and go to "System Expert > Hall A > Magnets > Big Bite (new version)".
  5. In the GUI, in the SETPOINT field, type the desired current value and press the enter/return key.
  6. Confirm that the READBACK field reports desired current value
  7. Make a follow-up HALOG post indicating that the BB magnet ramp is complete. Include a screenshot of the GUI in the post.

Ramp Magnet to 0 before powering off. Always have magnet off when switching polarity. Nice explanation is here

Hall Transition Procedures

REMINDER: The Run Coordinator will work with the Hall A Work Coordinator (Jessie Butler) to implement a hall transition.

The Super BigBite (SBS) magnet produces large fringe magnetic fields. When a radiation survey is performed, the RadCon team will use devices that can be compromised by magnetic fields, yielding incorrect readings. In magnetic fields stronger than 30 Gauss, ferrous materials like steel hand tools and other devices will experience a magnetic force and have the potential to become projectiles which can damage equipment and possibly cause serious injury to personnel. To mitigate this risk, Ops/Accelerator will use their procedure to ramp the SBS magnet and the corrector magnets down to zero current prior to personnel entering the hall under controlled access. We will additionally ramp the BigBite magnetic field down to zero current when personnel need to enter the Hall.

Procedure for Hall Transition to Controlled Access

The steps to ramp the BigBite magnet to zero field are as follows:

  1. Obtain permission from the Run Coordinator to ramp the BB magnetic field down to zero.
  2. Call MCC to confirm that they will ramp down the SBS magnet and communicate that you will ramp down the BB magnet.
  3. Make a post in the HALOG indicating that the BigBite magnetic field will be ramped to zero.
  4. Follow the directions above to ramp the BB field down to zero current.
  5. Make a follow-up HALOG post indicating that the BB magnet is at zero current. Include a screenshot of the GUI in the post.

Procedure for Hall Transition to Power/Beam Permit

To prepare the hall to receive beam and take data, we need to ramp the SBS and BB magnets to their full nominal currents. For SBS, the current is 2100 A. For BB, the current is 700 A. Ops/Accelerator will ramp up the SBS magnet and corrector magnets using their procedures. We ramp up the BigBite magnet with the following steps:

  1. Obtain permission from the Run Coordinator to ramp the BB magnetic field to its nominal current.
  2. Call MCC to confirm that they will ramp up the SBS and corrector magnets and communicate that you will ramp up the BB magnet.
  3. Make a post in the HALOG indicating that the BigBite magnetic field will be ramped to full field.
  4. Follow the directions above to ramp the BB magnet to 700 A.
  5. Make a follow-up HALOG post indicating that the BB magnet is at full current. Include a screenshot of the GUI in the post.

Making a Halog entry

You can either

  1. From a terminal on any of the desktop computers type "halog" and fill out the fields. You can include screenshots and attachments if useful.
  2. Navigate to the HALOG click "Add Content". Fill out the fields as necessary and submit it. This requires you to log in with you CUE username and password. Don't forget to log out.

Alarms

There are a number of alarms that shifters must respond to in Hall A in both the Counting House and the Hall. For information on how to respond to various alarms, click the links below

  • The Fire alarm. If this sounds and danger is imminent, immediately leave the building, call 911 and notify the guard shack 757-269-5822. If danger is not imminent, call for beam off and notify the crew chief, set the experiment in a safe state (if the target is moving for example, terminate the move), leave the building as soon as possible and notify the guard shack 757-269-5822.
  • VESDA (Very Early Smoke Detection Alarms). These are used to detect fire/smoke in the experimental hall. If they go off, check the hall camera for fire and notify the crew chief and the RC.
  • Cryo Target Alarm Handler The EPICS-based Target Alarm Handler is used to service only alarms specific to critical target function such as target movement, temperature and vacuum level. It is managed by Hall A staff (Silviu Covrig, Greg Smith, Jian-Ping Chen) and can only be changed upon request.
  • HallA Alarm Handler This EPICS-based Target Alarm Handler has the same underlying mechanics as the target alarm handler and is typically used to service critical hardware like spectrometer currents. It is managed by Hall A staff (Bob Michaels) and can only be changed upon request. To open, go to computer hadesk9 and login as adaq. Then 'ssh aslow@hadesk9' and then 'go_genrp'
  • Python Alarm Handler (Not utilized so far by GMn -- DO NOT USE) The Python-based alarm handler is administered by experts in the experiment (David Flay, Don Jones) and can be edited at our convenience to include or exclude variables and change thresholds etc.

Filling out BTA

  • Login to the BTA webpage with your CUE username. Notice the Help menu.
  • Move the mouse to "File". Don't click. Let mouse hover.
  • A dialog comes up. Select "Open Timesheet".
  • Select the Hall (A), and the shift, etc.
  • As a convenience feature, you may "Load from Epics" (button at bottom right) which automatically fills the fields. Note, the shift worker must click the little pencil-looking thing in the Edit column and MUST click the check mark that appears..
  • Note, automated numbers from EPICS works ONLY FOR PRODUCTION RUNNING with the SBS DAQ, i.e. it does not account for beam used by a variety of other tasks, such as energy measurements or HARP scans. The experimenters have final say over the numbers; you can edit the fields when you click Edit. (And don't forget to click the check mark whether you have edited or not; you MUST acknowledge.) Common sense: If the beam is used by the experiment it is ``ABU.
  • Submit on the "Signatures" tab.

Beam Line

Target

LHRS Information

LHRS for SBS

BigBite Spectrometer (BB)

GEMs

All GEM procedures and trouble shooting are located in this document GEM HOW TO

If the GEMs are causing the DAQ issues shift crews can reset everything through the GUI. From the a-onl@aonlX machine do:

gosbs
GEM_resets.sh
  • GEM HV is controlled through the main HV GUI explained here
  • Shift crews can check gas flow here:
  • Shift crews should check the gas pressures by first logging into a-onl@aonlX (X=2, 3, 4). Then
gosbs
CheckGEMGas.sh
  • On the aonl1 machine, to check the gas status, please do through following steps:
gosbs
module load epics
CheckGEMGas.sh

GRINCH

Timing Hodoscope

BBCAL: Calorimeter wiki

Optics

  • Straight-track optics running:
  1. Ask MCC to turn OFF SBS magnet and correctors. Ramp down the BB magnet in accordance with our procedures.
  2. Adjust target holding field for the BB and SBS magnets being off.
  3. Insert carbon foil target
  4. Load BBCal HV for PMTs for 0 field: [35]
  5. Set trigger for BB singles, and set threshold to 1.5-2 GeV for BBCal. Request 5uA beam and verify the rates. Increase the beam current (probably 10uA-ish) to keep DAQ around 3kHz and check the data. Continue for min 1.2M (1.5-1.8M ideal) events. If possible, go to 20 uA beam current.
  6. Update the DB for 0-field running:
In $SBS_REPLAY/DB/db_bb.dat, make an entry for the 0-field run as
bb.frontconstraintwidth_x = 0.75
bb.frontconstraintwidth_y = 0.25 
In $SBS_REPLAY/DB/20220701/db_bb.gem.dat, make an entry for the 0-field settings as:
bb.gem.useopticsconstraint = 0
bb.gem.useslopeconstraint = 0
bb.gem.useforwardopticsconstraint = 0


  • Multi-foil optics running:
  1. Ask MCC to turn ON the SBS magnet and correctors. Set BB magnet ON.
  2. Adjust target holding field for BB and SBS magnets being on.
  3. Insert the carbon optics target
  4. Load BBCal HV for PMTs with magnets ON: [36]
  5. Keeping thresholds from the straight-track optics running, ask for 5uA beam current to check rates. Increase current keeping the trigger ~3kHz and collect 17M events.
  6. Update the DB for field ON running:
In $SBS_REPLAY/DB/db_bb.dat, make an entry for the field ON run as:
bb.frontconstraintwidth_x = 0.15
bb.frontconstraintwidth_y = 0.11 
In $SBS_REPLAY/DB/20220701/db_bb.gem.dat, make an entry for the field ON settings as:
bb.gem.useopticsconstraint = 1
bb.gem.useslopeconstraint = 1
bb.gem.useforwardopticsconstraint = 1

Super BigBite Spectrometer (SBS)

SBS GEMs

HCAL: Calorimeter wiki

Experts: Jiwan Poudel (jpoudel@jlab.org), Sebastian Seeds (sseeds@jlab.org)